# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ 2 A classical discretization on a Shishkin mesh
We denote the $L^{2}(\Omega)$ norm by $\|.\|$ and the energy norm by
@@
where $H^{1}(\Omega)$ is the usual Sobolev space of functions whose first-order partial derivatives lie in $L^{2}(\Omega)$. We discretize (1.3) by means of the standard finite element method on a special rectangular mesh. Let $\mathcal{T}_{h}$${\{\tau_{i}^{h}\}}_{i=1}^{N}$ be a family of edge-to-edge rectangle’s of $\Omega$ with the parameter $h$.
$S_{h}=\{\mathcal{X}\in\mathcal{C}^{0}(\Omega),\mathcal{X}=0\;\text{on}\;\partial\Omega,\mathcal{X}|_{\tau_{i}^{h}}\;\text{linear in x and y}.\}$
The standard finite element discretization of (1.3) is:
Find $u_{h}\in S_{h}$ such that
$B_{h}(u_{h},v_{h})=(f,v_{h})\quad\forall\;v_{h}\in S_{h},$ (2.2)
where $B_{h}(.,.)$ is the bilinear form defined as
$B_{h}(u_{h},v_{h})=(\varepsilon\nabla u_{h},\nabla v_{h})+(\textbf{b}.\nabla u_{h},v_{h})+(cu_{h},v_{h})\quad\forall\;u_{h},v_{h}\in S_{h}.$ (2.3)
To effectively address the layering phenomenon inherent in the problem’s solution, we employ a layer-adapted mesh of Shishkin type, constructed independently along both the $x$ and $y$ spatial directions. For an even positive integer $N\geq 4$, the crucial mesh transition parameters in the $x$ and $y$ directions are denoted by $\lambda_{x}$ and $\lambda_{y}$, respectively. Note that these newly defined transition parameters are one of the novel contributions of this article. They are defined as,
$\lambda_{x}=min\left(\frac{2\varepsilon}{\alpha}\log\left(\frac{1}{\varepsilon}\right),\frac{1}{2}\right),\quad\lambda_{y}=min\left(2\sqrt{\frac{\varepsilon}{\beta}}\log\left(\frac{1}{\varepsilon^{3/2}}\right),\frac{1}{4}\right),$
@@
$x_{j}=\begin{cases}\frac{2\lambda_{x}}{N}i,&i=0,1,\dots,N/2,\\ \lambda_{x}+\frac{2}{N}(1-\lambda_{x})(i-\frac{N}{2}),&i=N/2+1,\dots,N.\end{cases}$ (2.4)
@@
In the $x$-direction, the interval $[0,1]$ is bisected into two sub-intervals $0,\lambda_{x}]$ and $\lambda_{x},1]$. Each of these sub-intervals is further partitioned such that its closure contains $N/2+1$ points, denoted as $X_{K}^{N}$. A symmetric partitioning is implicitly applied to [1,0][-1,0] to cover the full $x$ range. For the $y$-direction, the interval [1,1][-1,1] is divided into three sub-intervals: $1,-1+\lambda_{y}]$, $1+\lambda_{y},1-\lambda_{y}]$ and $1-\lambda_{y},1]$. These sub-intervals are then uniformly discretized. Specifically, their closures are partitioned into $N/4+1$, $N/2+1$, and $N/4+1$ points, respectively, forming the set $Y_{J}^{N}$. The global mesh points $\Omega^{N}$ are defined as the tensor product of these one-dimensional partitions:
@@
$h_{x}$ $=2\lambda_{x}N^{-1},\quad\quad H_{x}=2(1-\lambda_{x})N^{-1}$ (2.6)
$h_{y}$ $=4\lambda_{y}N^{-1},\quad\quad H_{y}=4(1-\lambda_{y})N^{-1}$ (2.7)
Here, $h_{x}$, $h_{y}$ denote the mesh widths within the layer regions in the $x$ and $y$ directions, respectively, while $H_{x}$, $H_{y}$ represent the mesh widths outside these regions. Due to the symmetry of the problem with respect to the $y$-axis, we analyze our problem in the reduced domain $\Omega=[0,1]\times[-1,1]$[7]. Within this domain, we perform a decomposition into four non-overlapping subregions, tailored to distinguish between areas of expected smooth behavior and those containing layers. This partition is defined as:
$\Omega_{c}=[\lambda_{x},1]\times[-1+\lambda_{y},1-\lambda_{y}],\quad\Omega_{f,x}=[0,\lambda_{x}]\times[-1+\lambda_{y},1-\lambda_{y}]$
$\Omega_{f,y}=([-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1])\times[\lambda_{x},1],\quad\Omega_{f,xy}=[0,\lambda_{x}]\times([-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1])$
3 Estimates for Discrete Green’s function
This section is devoted to finding out some important estimates for the discrete Green’s function corresponding to the differential operator in (1.1). These estimates play a fundamental role in deriving the pointwise error estimates. We derive the estimates of the discrete Green’s function in $L^{2}$ and energy norms. We note here that the Green’s function, or the discrete Green’s function, contains information about the source point. Therefore, these estimates play a crucial role in determining whether we can obtain an $\varepsilon$-independent estimate or not. We shall see that if the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then we obtain an $\varepsilon$-independent estimate for the Green’s function in energy and $L^{2}$ norms, and we can obtain a parameter-uniform $L^{\infty}$ norm error estimate.
Lemma 3.1 When the source point $(x_{m},y_{n})\in\Omega_{c}$, then $\|G_{mn}\|_{1,\varepsilon}\leq CN.$
Proof. From the coercivity of the bilinear form we have,
@@ 3 Estimates for Discrete Green’s function
$\leq CN^{2}\|G_{mn}\|_{L^{1}(R_{mn})}\quad\quad\quad(\text{by inverse estimate})$
@@
$\|G_{mn}\|_{{1,\varepsilon}}$ $\leq CN.$
Corollary 1 When $(x_{m},y_{n})\in\Omega_{c}$ then $\|G_{mn}\|\leq CN.$
Lemma 3.2 When $(x_{m},y_{n})\in\Omega_{f,x}$ then $\|G_{mn}\|_{1,\varepsilon}\leq CN^{1/2}\log^{1/2}(1/\varepsilon).$
Proof. From the coercivity of the bilinear form we have,
@@
$\leq CN^{1/2}\log^{1/2}(1/\varepsilon)\|G_{mn}\|_{1,\varepsilon}.$
$\|G_{mn}\|_{1,\varepsilon}$ $\leq CN^{1/2}\log^{1/2}(1/\varepsilon).$
Corollary 2 When $(x_{m},y_{n})\in\Omega_{f,x}$ then $\|G_{mn}\|\leq C\varepsilon^{-1/2}\log^{-1/2}(1/\varepsilon).$
@@
Remark 1 Due to the $\varepsilon$-dependent estimate for $\|G_{mn}\|$ when $(x_{m},y_{n})\in\Omega_{f,y}\cup\Omega_{f,xy}$ it is not possible to obtain $\varepsilon$-independent error estimate. Note that when $(x_{m},y_{n})\in\Omega_{f,y}$ we are able to obtain $\|G_{mn}\|\leq C\varepsilon^{-1/4}\log^{-1/2}(1/\varepsilon)$ and when $(x_{m},y_{n})\in\Omega_{f,xy}$ then $\|G_{mn}\|\leq C\varepsilon^{-1/4}$. We are not able to obtain a parameter uniform estimate for $\|G_{mn}\|_{1,\varepsilon}$ when the source point $(x_{m},y_{n})\in\Omega_{f,y}\cup\Omega_{f,xy}$, which implies the $\varepsilon$-dependent error estimate in $L^{\infty}$ norm, which is also reflected in the numerical examples.
Theorem 2
@@
[7] Assume that $K\in\mathcal{J}_{N}$ is a mesh rectangle with sides that are perpendicular to the coordinate axes. Assume that $w\in H^{3}(K)$ and that $w^{I}$ is its nodal bilinear interpolate on $K$. Hence, we obtain for any bilinear function $v^{N}$ defined on $K$
@@
Lemma 4.1 If the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then $\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega}\leq\frac{C(\log(1/\varepsilon^{3/2}))^{3}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$.
$\varepsilon(\nabla(R-IR),\nabla G_{mn})$ $=\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,xy}}+\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,x}}$
$+\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,y}}+\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}$ (3.1)
Now, consider the first term of the R.H.S. of (3.1) in the $x$-direction $\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,xy}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{9/4}(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.2)
Now, consider the first term of the R.H.S. of (3.1) in the $y$-direction $\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,xy}}$. We obtain,
$\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,xy}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,xy}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;N$
$\leq C\frac{\varepsilon^{13/4}(\log(1/\varepsilon^{3/2}))^{3}}{N}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$ (3.3)
Now, consider the second term of the R.H.S. of (3.1) in the $x$-direction $\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,x}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{1}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\;N$
$\leq C\frac{\varepsilon(\log(1/\varepsilon))^{1/2}}{N}\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N}.$ (3.4)
Now, consider the second term of the R.H.S. of (3.1) in the $y$-direction $\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,x}}$. We obtain,
$\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,x}}$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,x}}}$

### v2

@@ 2 A classical discretization on a Shishkin mesh
We denote the $L^{2}(\Omega)$ norm by $\|.\|$ and the energy norm by $\|.\|_{1,\varepsilon}$ and is defined by
@@
where $H^{1}(\Omega)$ is the usual Sobolev space of functions whose first-order weak derivatives lie in $L^{2}(\Omega)$. We discretize (1.3) by means of the standard finite element method on a special rectangular mesh. Let $\mathcal{T}_{h}$${\{\tau_{i}^{h}\}}_{i=1}^{N}$ be a family of edge-to-edge rectangle’s whose union is $\bar{\Omega}$. Define
$S_{h}=\{\mathcal{X}\in\mathcal{C}^{0}(\bar{\Omega}):\mathcal{X}=0\;\text{on}\;\partial\Omega,\mathcal{X}|_{\tau_{i}^{h}}\;\text{linear in $x$ and $y$}\},$
where $\mathcal{C}^{0}(\bar{\Omega})$ is space of continuous functions up-to boundary of $\Omega$. The standard finite element discretization of (1.3) is to find $u_{h}\in S_{h}$ such that
$B(u_{h},v_{h})=(f,v_{h})\quad\forall\;v_{h}\in S_{h}.$ (2.2)
To effectively address the layering phenomenon inherent in the problem’s solution, we employ a layer-adapted mesh of Shishkin type, constructed independently along both the $x$ and $y$ spatial directions. For an even positive integer $N\geq 4$, the crucial mesh transition parameters in the $x$ and $y$ directions are denoted by $\lambda_{x}$ and $\lambda_{y}$, respectively. Note that these newly defined transition parameters are one of the novel contributions to this article. They are defined as:
$\lambda_{x}=min\left(\frac{2\varepsilon}{\alpha}\log\left(\frac{1}{\varepsilon}\right),\frac{1}{2}\right),\quad\lambda_{y}=min\left(2\sqrt{\frac{\varepsilon}{\beta}}\log\left(\frac{1}{\varepsilon^{3/2}}\right),\frac{1}{4}\right).$
@@
$x_{i}=\begin{cases}\frac{2\lambda_{x}}{N}i,&i=0,1,\dots,N/2,\\ \lambda_{x}+\frac{2}{N}(1-\lambda_{x})(i-\frac{N}{2}),&i=N/2+1,\dots,N,\end{cases}$ (2.3)
@@
In the $x$-direction, the interval $[0,1]$ is bisected into two sub-intervals $0,\lambda_{x}]$ and $\lambda_{x},1]$. Each of these sub-intervals is further partitioned such that its closure contains $N/2+1$ points, denoted as $X_{K}^{N}$. A symmetric partitioning is implicitly applied to [1,0][-1,0] to cover the full $x$ range. For the $y$-direction, the interval [1,1][-1,1] is divided into three sub-intervals: $1,-1+\lambda_{y}]$, $1+\lambda_{y},1-\lambda_{y}]$ and $1-\lambda_{y},1]$. These sub-intervals are uniformly discretized. Specifically, their closures are partitioned into $N/4+1$, $N/2+1$, and $N/4+1$ points, respectively, forming the set $Y_{J}^{N}$. The collection of global mesh points is denoted by $\Omega^{N}$ and defined as the tensor product of these one-dimensional partitions:
@@
$h_{x}$ $=2\lambda_{x}N^{-1},\quad\quad H_{x}=2(1-\lambda_{x})N^{-1},$
$h_{y}$ $=4\lambda_{y}N^{-1},\quad\quad H_{y}=4(1-\lambda_{y})N^{-1}.$
Here, $h_{x}$, $h_{y}$ denote the mesh widths within the layer regions in the $x$ and $y$ directions, respectively, while $H_{x}$, $H_{y}$ represent the mesh widths outside these regions. Due to the symmetry of the problem with respect to the $y$-axis, we analyze our problem in the reduced domain $\Omega=[0,1]\times[-1,1]$. Within this domain, we perform a decomposition into four non-overlapping sub-regions, tailored to distinguish between areas of expected smooth behavior and those containing layers. This partition is defined as:
$\Omega_{c}=[\lambda_{x},1]\times[-1+\lambda_{y},1-\lambda_{y}],\quad\Omega_{f,x}=[0,\lambda_{x}]\times[-1+\lambda_{y},1-\lambda_{y}]$,
$\Omega_{f,y}=([-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1])\times[\lambda_{x},1],\quad\Omega_{f,xy}=[0,\lambda_{x}]\times\left([-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1]\right)$.
This discretization is followed from [21].
3 Estimates for discrete Green’s function and some auxiliary estimates
This section is devoted to finding out some important estimates for the discrete Green’s function corresponding to the differential operator in (1.2) and some key auxiliary estimates. These estimates play a fundamental role in deriving the pointwise error estimates. We derive the estimates for the discrete Green’s function in $L^{2}$-norm and energy norm. We note that here the discrete Green’s function contains information about the source point. Therefore, these estimates are essential for identifying regions where an $\varepsilon$-independent error estimate is feasible. We shall see that if the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then we obtain an $\varepsilon$-independent estimate for the discrete Green’s function in energy norm and $L^{2}$-norm, and we can expect to obtain a parameter-uniform $L^{\infty}$-norm error estimate. In this article, we obtain convergent error estimates specifically for the $x$-layer region, denoted as $\Omega_{f,x}$. However, for the coercive region, we can only establish an upper bound for the error that is independent of $\varepsilon$. This observation is consistent with the findings reported in [23] regarding the coercive region, further validating our analyses.
Lemma 3.1
When the source point $(x_{m},y_{n})\in\Omega_{c}$, then $\|G_{mn}\|_{1,\varepsilon}\leq CN.$
Proof. From the coercivity of the bilinear form, we have
@@ 3 Estimates for discrete Green’s function and some auxiliary estimates
$\leq CN^{2}\|G_{mn}\|_{L^{1}(R_{mn})},\hskip 56.9055pt(\text{by inverse estimate})$
@@
$\implies\;\|G_{mn}\|_{{1,\varepsilon}}$ $\leq CN.$
Corollary 3.1.1
When $(x_{m},y_{n})\in\Omega_{c}$, then $\|G_{mn}\|\leq CN.$
Lemma 3.2
When $(x_{m},y_{n})\in\Omega_{f,x}$, then $\|G_{mn}\|_{1,\varepsilon}\leq CN^{1/2}\log^{1/2}(1/\varepsilon).$
Proof. From the coercivity of the bilinear form, we have
@@
$\leq CN^{1/2}\log^{1/2}(1/\varepsilon)\|G_{mn}\|_{1,\varepsilon}$
$\implies\;\|G_{mn}\|_{1,\varepsilon}$ $\leq CN^{1/2}\log^{1/2}(1/\varepsilon).$
Corollary 3.2.1
When $(x_{m},y_{n})\in\Omega_{f,x}$, then $\|G_{mn}\|\leq CN^{1/2}\log^{1/2}(1/\varepsilon).$
Remark 3.2.1
@@
Note that when $(x_{m},y_{n})\in\Omega_{f,y}$ we are able to obtain $\|G_{mn}\|\leq C\varepsilon^{-1/4}\log^{-1/2}(1/\varepsilon)$ and when $(x_{m},y_{n})\in\Omega_{f,xy}$, then $\|G_{mn}\|\leq C\varepsilon^{-1/4}$. Due to the $\varepsilon$-dependent estimate for $\|G_{mn}\|$ when $(x_{m},y_{n})\in\Omega_{f,y}\cup\Omega_{f,xy}$ it is not possible to obtain $\varepsilon$-independent error estimate. Therefore, it is expected that we do not obtain parameter-independent error estimation when the source point $(x_{m},y_{n})\in\Omega_{f,y}\cup\Omega_{f,xy}$. This is reflected in the numerical example.
Lemma 3.3
If $(x,y)\in\Omega_{c}\cup\Omega_{f,y}$, then following estimates are holds
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}\leq C\varepsilon^{\alpha},\quad\quad\quad\;\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{c})}\leq C\varepsilon^{\alpha}$.
Proof. If $(x,y)\in\Omega_{f,y}$, then $x\in[\lambda_{x},1]$, $y\in[-1,-1+\lambda_{y}]\cup[1-\lambda_{y},1]$ and when $(x,y)\in\Omega_{c}$, then $x\in[\lambda_{x},1]$, $y\in[-1+\lambda_{y},1-\lambda_{y}]$.
In the both cases when $x\in[\lambda_{x},1]$, we obtain
$\implies e^{\frac{-\alpha x}{\varepsilon}}\leq e^{\frac{-\alpha\lambda_{x}}{\varepsilon}}=e^{-\frac{\alpha}{\varepsilon}\varepsilon\log\left(\frac{1}{\varepsilon}\right)}=e^{-\alpha\log\left(\frac{1}{\varepsilon}\right)}=e^{\log\left(\frac{1}{\varepsilon^{-\alpha}}\right)}=\varepsilon^{\alpha}.$
This completes the proof.
Corollary 3.3.1
If $(x,y)\in\Omega_{c}\cup\Omega_{f,x}$, then following estimates are holds
$\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,x})}=1\quad\text{and}\quad\|e^{-\frac{\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,xy})}=1.$
Lemma 3.4
If $(x,y)\in\Omega_{c}\cup\Omega_{f,x}$, then following estimates are holds
$\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}\leq C\varepsilon^{\frac{3\beta}{2}},$
$\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\leq C\varepsilon^{\frac{3\beta}{2}}.$
Proof. If $(x,y)\in\Omega_{c}$, then $x\in[\lambda_{x},1]$, $y\in[-1+\lambda_{y},1-\lambda_{y}]$, we have
$\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{c})}$ $=\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}$
$\leq\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}+\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})},$
and when $(x,y)\in\Omega_{f,x}$, then $x\in[0,\lambda_{x}]$ and $y\in[-1+\lambda_{y},1-\lambda_{y}]$. We have
$\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}$ $=\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}$
$\leq\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}+\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}.$
For the both cases we need to calculate $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}$. As
$y\in[-1+\lambda_{y},1-\lambda_{y}]\implies-1+\lambda_{y}\leq y\leq 1-\lambda_{y}$
First we take $1+y\geq\lambda_{y},$ then
$-\frac{\beta}{\sqrt{\varepsilon}}(1+y)\leq-\frac{\beta}{\sqrt{\varepsilon}}\lambda_{y}\implies e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\leq e^{-\frac{\beta}{\sqrt{\varepsilon}}\lambda_{y}}$
$\implies e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\leq e^{-\frac{\beta}{\sqrt{\varepsilon}}\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{-\beta\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{\log\left(\frac{1}{\varepsilon^{\frac{-3\beta}{2}}}\right)}=\varepsilon^{\frac{3\beta}{2}},$
and if $y\leq 1-\lambda_{y},$ then
$-\frac{\beta}{\sqrt{\varepsilon}}(1-y)\leq-\frac{\beta}{\sqrt{\varepsilon}}\lambda_{y}\implies e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\leq e^{-\frac{\beta}{\sqrt{\varepsilon}}\lambda_{y}}$
$\implies e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\leq e^{-\frac{\beta}{\sqrt{\varepsilon}}\sqrt{\varepsilon}\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{-\beta\log\left(\frac{1}{\varepsilon^{3/2}}\right)}=e^{\log\left(\frac{1}{\varepsilon^{\frac{-3\beta}{2}}}\right)}=\varepsilon^{\frac{3\beta}{2}}.$
This completes the proof.
Corollary 3.4.1
If $(x,y)\in\Omega_{f,y}\cup\Omega_{f,xy}$, then following estimates are holds
$\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,y})}=1,$ and $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,xy})}=1.$
4 Estimates related to diffusion terms
In this section, we aim to bound the diffusion portion of the bilinear form B(.,.) with its arguments parts of interpolation error (uIu)(u-Iu) and the discrete Green’s function $G_{mn}.$ We start by presenting the following auxiliary results.
Theorem 4.1
@@ 4 Estimates related to diffusion terms
[21] Assume that $K\in\mathcal{J}_{N}$ is a mesh rectangle with sides that are perpendicular to the coordinate axes. Assume that $u\in H^{3}(K)$ and that $Iu$ is its nodal bilinear interpolate on $K$. Hence, we obtain for any bilinear function $v^{N}$ defined on $K$, then
@@
Lemma 4.2
If source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then $|\varepsilon(\nabla(R-IR),\nabla G_{mn})|\leq\frac{C(\log(1/\varepsilon^{3/2}))^{3}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{f,x}$ and decompose the domain in four parts, we have
$|\varepsilon(\nabla(R-IR),\nabla G_{mn})|$ $\leq|\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,xy}}|+|\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,x}}|$
$+|\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{f,y}}|+|\varepsilon(\nabla(R-IR),\nabla G_{mn})_{\Omega_{c}}|.$ (4.1)
Now, estimate the first term of the R.H.S. of (4.1) in the $x$-direction is $|\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1 and furthermore, we obtain
$|\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,xy}}|$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,xy}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\varepsilon^{9/4}(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}.$ (4.2)
Now, estimate the first term of the R.H.S. of (4.1) in the $y$-direction is $|\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,xy}}|$. By using Theorem 1.2, Theorem 4.1 and furthermore, we obtain
$|\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,xy}}|$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,xy}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,xy}}}$
@@
$\leq C\sqrt{\varepsilon}\left(\frac{\varepsilon\log(1/\varepsilon)}{N}\right)^{2}\sqrt{\varepsilon\log(1/\varepsilon)}\sqrt{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}\;\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\frac{\varepsilon^{13/4}(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N^{3/2}}.$ (4.3)
Now, estimate the second term of the R.H.S. of (4.1) in the $x$-direction is $|\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,x}}|$. By using 1.2, Theorem 4.1 and furthermore, we obtain
$|\varepsilon((R-IR)_{x},G_{mn,x})_{\Omega_{f,x}}|$ $\leq C\sqrt{\varepsilon}\;h_{y}^{2}\;\|{R}_{xyy}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,x}\|_{{\Omega_{f,x}}}$
@@
$\leq C\frac{\varepsilon\log(1/\varepsilon)}{N^{3/2}}\leq C\frac{\log(1/\varepsilon)^{1/2}}{N^{3/2}}.$ (4.4)
Now, estimate the second term of the R.H.S. of (4.1) in the $y$-direction is $|\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1 and furthermore, we obtain
$|\varepsilon((R-IR)_{y},G_{mn,y})_{\Omega_{f,x}}|$ $\leq C\sqrt{\varepsilon}\;h_{x}^{2}\;\|{F_{1}}_{yxx}\|_{\Omega_{f,x}}\;\sqrt{\varepsilon}\|G_{mn,y}\|_{{\Omega_{f,x}}}$